the two congruent angles of an isoseles triangle measure (2x-20)degrees each. The third angle measures x degrees. What is the measure of each angle?
Found 2 solutions by rfer, checkley77:
Answer by rfer(16322)

(2x-20)+(2x-20)+x=180
2x-20+2x-20+x=180
5x=220
x=44 degrees
2x-20=68 degrees

Answer by checkley77(12844)

2(2x-20)+x=180
4x-40+x=180
5x=180+40
5x=220
x=220/5
x=44 degrees for the third angle.
2*44-20=88-20=68 degrees for the two concruent angles.
Proof:
2*68=44=180
136+44=180
180=180